# ReHub: Linear Complexity Graph Transformers with Adaptive Hub-Spoke Reassignment

### v1

@@ Abstract
We present ReHub, a novel graph transformer architecture that achieves linear complexity through an efficient reassignment technique between nodes and virtual nodes. Graph transformers have become increasingly important in graph learning for their ability to utilize long-range node communication explicitly, addressing limitations such as oversmoothing and oversquashing found in message-passing graph networks. However, their dense attention mechanism scales quadratically with the number of nodes, limiting their applicability to large-scale graphs. ReHub draws inspiration from the airline industry’s hub-and-spoke model, where flights are assigned to optimize operational efficiency. In our approach, graph nodes (spokes) are dynamically reassigned to a fixed number of virtual nodes (hubs) at each model layer. Recent work, Neural Atoms (Li et al., 2024), has demonstrated impressive and consistent improvements over GNN baselines by utilizing such virtual nodes; their findings suggest that the number of hubs strongly influences performance. However, increasing the number of hubs typically raises complexity, requiring a trade-off to maintain linear complexity. Our key insight is that each node only needs to interact with a small subset of hubs to achieve linear complexity, even when the total number of hubs is large. To leverage all hubs without incurring additional computational costs, we propose a simple yet effective adaptive reassignment technique based on hub-hub similarity scores, eliminating the need for expensive node-hub computations. Our experiments on long-range graph benchmarks indicate a consistent improvement in results over the base method, Neural Atoms, while maintaining a linear complexity instead of $O(n^{3/2})$. Remarkably, our sparse model achieves performance on par with its non-sparse counterpart. Furthermore, ReHub outperforms competitive baselines and consistently ranks among the top performers across various benchmarks.
@@ 1 Introduction
In this work, we introduce ReHub, a novel graph transformer architecture that achieves linear complexity by dynamically reassigning graph nodes to virtual nodes. We are inspired by complex systems where efficient connectivity and adaptability are crucial for optimal performance. A pertinent example is the airline industry, where flights are dynamically assigned to a limited number of major airports (hubs) to optimize operational efficiency. We identify the graph nodes with spokes and the virtual nodes with hubs. The key insight of our approach stems from noting that the transformer’s complexity is driven by spoke-hub attention. Therefore, it is not necessary to reduce the total number of hubs to achieve linear complexity; instead, using a fixed, small number of connected hubs per spoke is sufficient. To effectively utilize all hubs without increasing computational cost, we introduce a simple yet efficient adaptive reassignment mechanism. In order to so and yet avoid the costly computation the entire set of spoke-hub interactions, our reassignment mechanism is based on hub-hub similarity scores, which are cheap to compute.
@@ 3.1 Pipeline overview
Figure 1: Illustration of ReHub architecture. (1) Overview of the different steps in the architecture. $\mathcal{A}$ is the input (spoke) graph’s adjacency matrix; spoke features $\bm{s}$; hub features $\bm{h}$; hub assignment $\bm{E}$; and $\bm{\Gamma}$ contains attention scores from the Hubs-Spokes attention. (2) Hubs initialization. Spokes are first clustered, then each cluster is aggregated to compute hub features. Finally, each spoke is assigned more hubs. (3) Hub (Re)Assignment. A spoke connected to $k$ hubs updates its hub assignment to the $k$ hubs closest to the hub most similar to it. (4) Illustration of connectivity between spokes and hubs. Information pass between spokes with an MPNN; while the interaction between hubs and spokes is performed via attention and only through available connections $E_{i_{s}i_{h}}$. The hubs pass information between themselves via full self-attention.
@@ 3.2 Notation
${\bm{E}}_{i_{s},i_{h}}=\begin{cases}1&\text{if spoke $i_{s}$ is connected to hub $i_{h}$}\\ 0&\text{otherwise}\end{cases}\quad\text{s.t.}\quad{\bm{E}}\,{\bm{1}}_{N_{h}}=k\cdot{\bm{1}}_{N_{s}},$ (1)
@@ 3.3 Initialization
Hub assignment.  We aim to connect each spoke with a total of $k$ hubs. Since each spoke $\bm{s}_{i_{s}}$, already has a single connection to a hub $\bm{h}_{i_{h}}$ as a result of the clustering, we assign the remaining $k-1$ hub connections for each spoke, by selecting the hubs closest to $\bm{h}_{i_{h}}$ in terms of feature similarity. The assignment procedure is detailed in Section 3.4. This step results in the spoke-to-hub initial assignment matrix ${\bm{E}}^{0}_{s\rightarrow h}$.
@@ 3.4 Long-range spoke update layer
(5) Hub (Re)Assignment:  While restricting each spoke to connect with $k$ hubs maintains efficiency, for the method to achieve its full potential, it should utilize all available hubs. We achieve this by keeping only $k$ connected hubs per spoke at each layer, while allowing each spoke to reassign $k-1$ of its hubs before proceeding to the next layer [^2] We note that, although we replace $k-1$ hubs while keeping the closest hub connected, in later layers, that closest hub may change. This flexibility prevents us from being overly constrained by the initial hub selection.. The reassignment is based on spoke-hub similarity. Selecting the hubs closest to each spoke, however, would require $N_{s}\times N_{h}$ computations. To avoid this, we leverage the distances between hubs, which are efficient to compute.
We then retain the hub most similar to each spoke from the sparse set of connected hubs, replacing the remaining $k-1$ hubs with those closest to it. This procedure is outlined in Algorithm 1. The matrix of distances between all hub features is denoted by $\bm{\Delta}$. Naturally, a hub is closest to itself and would be selected first. Additionally, we use $\bm{\Gamma}$, the attention score matrix, to identify, for each spoke, the most similar hub to which it is connected.
@@ 4 Experiments
Datasets.  We evaluate ReHub on (1) long-range communication ability and (2) large graphs to verify memory efficiency. For long-range communication we evaluate ReHub on the long-range graph benchmark (LRGB) which is is widely used to evaluate methods which aim at overcoming issues such as oversmoothing and oversquashing. LRGB comprises five datasets. Two of the datasets are image-based graph datasets: PascalVOC-SP and COCO-SP which contain superpixel graphs of the well known image segmentation datasets PascalVOC and COCO. The latter three datasets are molecular datasets: Peptides-Func, Peptides-Struct and PCQM-Contact, which require the prediction of molecular interactions and properties that require global aggregation of information. For evaluation on large graphs we show results on graph datasets of citation networks: OGBN-Arxiv and Coauthor Physics which include about 170K and 30K nodes respectively, with the task of node class prediction. Additionally, we evaluate peak memory consumption on a custom dataset of large random regular graphs (Steger & Wormald, 1999; Kim & Vu, 2003) of gradually increasing sizes from 10K to 700K nodes. Additional statistics about the datasets is available in Appendix A.1
Metrics and evaluation.  We evaluate our models using several metrics: Average Precision (AP), Mean Absolute Error (MAE), Mean Reciprocal Rank (MRR), F1 Score, and Accuracy. These are standard metrics and we refer to Dwivedi et al. (2022) for more details. For the evaluation of ReHub, we report the mean $\pm$ std over 5 runs, each with a different random seed.
@@ 4.3 Ablations and analyses
Long-range spoke update layer components.  As described in Section 3, our long-range spoke update layer is designed to handle long range communication while maintaining memory efficiency. The ablation provided in Table 4 highlights the contributions of the primary design choices. (1) Initialization of hub features from spokes vs. learned hub features as parameters, where the latter is analogous to the initialization process of Neural Atoms. The initialization scheme, outlined in Section 3.3, is compared to a configuration where hub features are learned parameters, i.e, $\bm{h}_{i_{h}}^{0}={\bm{P}}_{i_{h}}\in\mathbb{R}^{d}$. Initializing the hubs from spokes significantly improves performance, while using learned hubs results in reduced performance compared to the GNN baseline. (2) The use of a fixed vs. dynamic number of hubs. For the fixed case, we set the number of hubs to 22, which is approximately the square root of the average number of nodes, $\sqrt{479.4}\approx 21.9$. This is compared to setting the number of hubs as $N_{h}=\sqrt{N_{s}}$ dynamically, according to graph size. (3) Reassigning spokes to hubs at each layer vs. keeping the same connections fixed across the layers, is shown to improve performance. (4) Including an encoding layer for the spokes prior to aggregation further improves performance.
Table 4: Ablation study. We measure the effect of various components of ReHub on top of a GatedGCN MPNN, using the PascalVOC-SP dataset. The number of hubs used per graph (#Hubs): for 22 it is a static amount and for $\sqrt{N_{s}}$ it is dynamic per graph size. Initial hubs (Hubs Init) can be set as learned parameters or initialized from the assigned spokes as described in 3.3 where we can add a feedforward layer on the spokes (Spokes Enc) before aggregation. Reassignment is as described in 3.4. We use $k=3$ for all runs.
@@
Hub utilization.  To gain insights into the reassignment dynamics, we measure the level of hub utilization. We define hub utilization $U$ as the number of hubs that have at least one spoke connected to them. Formally, $U=|\{i_{h}\mid{\bm{E}}_{:,i_{h}}\cdot{\bm{1}}_{N_{h}}\geq 1\}|$ where ${\bm{E}}_{:,i_{h}}$ represents the set of all spokes connected to hub $i_{h}$. The percentage of unused hubs is then given by $1-U/N_{h}$, which reflects the proportion of hubs that are not connected to any spoke. Figure 3 presents a cumulative graph illustrating the percentage of graphs with an unused hub percentage below a given threshold. Hub utilization is shown per layer for different configurations of hub ratio $r$ and connected hubs $k$, based on the validation split of the PascalVOC-SP dataset. The results consistently demonstrate that approximately 10% of hubs remain isolated in each graph, across layers, regardless of the number of hubs or the number of spokes per hub. This finding suggests that the network maintains robust information flow between spokes and hubs. We also present an additional metric based on the Bhattacharyya coefficient in Appendix A.4.
Figure 3: Hub utilization. Unused hubs distribution for varying $k$ and hub ratio $r$ evaluated on the validation split of PascalVOC-SP dataset. Left: $k=3$ with $r\in\{1,4\}$. Right: $k=5$ with $r\in\{1,4\}$.
@@ 5 Conclusion and future work
Limitations  While consistently improving computational efficiency, our spoke-hub communication maintained strong performance, though it did not establish a new state of the art. Additionally, our current design lacks inherent support for positional information available in geometric graphs. In future work, we aim to extend our method to support tasks requiring long-range communication on geometric graphs by incorporating positional information into the spoke-hub attention and reassignment mechanisms. We also plan to make the reassignment module learnable, and optimize it towards the prediction task to further boost accuracy. Finally, integrating ReHub into architectures like Exphormer to enable the reassignment of expander graph edges between layers presents a promising direction to further enhance long-range communication.
Acknowledgments
Or Litany is a Taub fellow and is supported by the Azrieli Foundation Early Career Faculty Fellowship.

### v2

@@ Abstract
We present ReHub, a novel graph transformer architecture that achieves linear complexity through an efficient reassignment technique between nodes and virtual nodes. Graph transformers have become increasingly important in graph learning for their ability to utilize long-range node communication explicitly, addressing limitations such as oversmoothing and oversquashing found in message-passing graph networks. However, their dense attention mechanism scales quadratically with the number of nodes, limiting their applicability to large-scale graphs. ReHub draws inspiration from the airline industry’s hub-and-spoke model, where flights are assigned to optimize operational efficiency. In our approach, graph nodes (spokes) are dynamically reassigned to a fixed number of virtual nodes (hubs) at each model layer. Recent work, Neural Atoms (Li et al., 2024), has demonstrated impressive and consistent improvements over GNN baselines by utilizing such virtual nodes; their findings suggest that the number of hubs strongly influences performance. However, increasing the number of hubs typically raises complexity, requiring a trade-off to maintain linear complexity. Our key insight is that each node only needs to interact with a small subset of hubs to achieve linear complexity, even when the total number of hubs is large. To leverage all hubs without incurring additional computational costs, we propose a simple yet effective adaptive reassignment technique based on hub-hub similarity scores, eliminating the need for expensive node-hub computations. Our experiments on long-range graph benchmarks indicate a consistent improvement in results over the base method, Neural Atoms, while maintaining a linear complexity instead of $O(n^{3/2})$. Remarkably, our sparse model achieves performance on par with its non-sparse counterpart. Furthermore, ReHub outperforms competitive baselines and consistently ranks among the top performers across various benchmarks. Code is available on our project webpage [^1] https://tomerborreda.github.io/rehub/.
@@ 1 Introduction
In this work, we introduce ReHub, a novel graph transformer architecture that achieves linear complexity by dynamically reassigning graph nodes to virtual nodes. We are inspired by complex systems where efficient connectivity and adaptability are crucial for optimal performance. A pertinent example is the airline industry, where flights are dynamically assigned to a limited number of major airports (hubs) to optimize operational efficiency. We identify the graph nodes with spokes and the virtual nodes with hubs. The key insight of our approach stems from noting that the transformer’s complexity is driven by spoke-hub attention. Therefore, it is not necessary to reduce the total number of hubs to achieve linear complexity; instead, using a fixed, small number of connected hubs per spoke is sufficient. To effectively utilize all hubs without increasing computational cost, we introduce a simple yet efficient adaptive reassignment mechanism. In order to do so, and yet avoid the costly computation of the entire set of spoke-hub interactions, our reassignment mechanism is based on hub-hub similarity scores, which are cheap to compute.
@@ 3.1 Pipeline overview
Figure 1: Illustration of ReHub architecture. (1) Overview of the different steps in the architecture. $\mathcal{A}$ is the input (spoke) graph’s adjacency matrix; spoke features $\bm{s}$; hub features $\bm{h}$; hub assignment $\bm{E}$; and $\bm{\Gamma}$ contains attention scores from the Hubs-Spokes attention. (2) Hubs initialization. Spokes are first clustered, then each cluster is aggregated to compute hub features. Finally, each spoke is assigned more hubs. (3) Hub (Re)Assignment. Each spoke selects the $k$ hubs closest to its most similar hub as its new assignment. (4) Illustration of connectivity between spokes and hubs. Information pass between spokes with an MPNN; while the interaction between hubs and spokes is performed via attention and only through available connections $E_{i_{s}i_{h}}$. The hubs pass information between themselves via full self-attention.
@@ 3.2 Notation
[…]
@@ 3.3 Initialization
Hub assignment.  We aim to connect each spoke with a total of $k$ hubs. Since each spoke $\bm{s}_{i_{s}}$, already has a single connection to a hub $\bm{h}_{i_{h}}$ as a result of the clustering, we assign the remaining k1k-1 hub connections for each spoke, by selecting the hubs closest to $\bm{h}_{i_{h}}$ in terms of feature similarity. The assignment procedure is detailed in Section 3.4. This step results in the spoke-to-hub initial assignment matrix ${\bm{E}}^{0}_{s\rightarrow h}$.
@@ 3.4 Long-range spoke update layer
(5) Hub (Re)Assignment:  While restricting each spoke to connect with $k$ hubs maintains efficiency, for the method to achieve its full potential, it should utilize all available hubs. We achieve this by keeping only $k$ connected hubs per spoke at each layer, while allowing each spoke to reassign k1k-1 of its hubs before proceeding to the next layer [^3] We note that, although we replace k1k-1 hubs while keeping the closest hub connected, in later layers, that closest hub may change. This flexibility prevents us from being overly constrained by the initial hub selection.. The reassignment is based on spoke-hub similarity. Selecting the hubs closest to each spoke, however, would require $N_{s}\times N_{h}$ computations. To avoid this, we leverage the distances between hubs, which are efficient to compute.
We then retain the hub most similar to each spoke from the sparse set of connected hubs, replacing the remaining k1k-1 hubs with those closest to it. This procedure is outlined in Algorithm 1. The matrix of distances between all hub features is denoted by $\bm{\Delta}$. Naturally, a hub is closest to itself and would be selected first. Additionally, we use $\bm{\Gamma}$, the attention score matrix, to identify, for each spoke, the most similar hub to which it is connected.
@@ 4 Experiments
Datasets.  Due to lack of large scale datasets that explicitly benchmark long-range capabilities we divide the evaluation to (1) long-range communication ability and (2) large graphs to verify memory efficiency. For long-range communication we evaluate ReHub on the long-range graph benchmark (LRGB) which is is widely used to evaluate methods which aim at overcoming issues such as oversmoothing and oversquashing. LRGB comprises five datasets. Two of the datasets are image-based graph datasets: PascalVOC-SP and COCO-SP which contain superpixel graphs of the well known image segmentation datasets PascalVOC and COCO. The latter three datasets are molecular datasets: Peptides-Func, Peptides-Struct and PCQM-Contact, which require the prediction of molecular interactions and properties that require global aggregation of information. For evaluation on large graphs we show results on graph datasets of citation networks: OGBN-Arxiv and Coauthor Physics which include about 170K and 30K nodes respectively, with the task of node class prediction. Additionally, we evaluate peak memory consumption on a custom dataset of large random regular graphs (Steger & Wormald, 1999; Kim & Vu, 2003) of gradually increasing sizes from 10K to 700K nodes. Additional statistics about the datasets is available in Appendix A.1
Metrics and evaluation.  We evaluate our models using several metrics: Average Precision (AP), Mean Absolute Error (MAE), Mean Reciprocal Rank (MRR), F1 Score, and Accuracy. These are standard metrics and we refer to Dwivedi et al. (2022) for more details. For the evaluation of ReHub, we report the mean $\pm$ std over 5 runs, each with a different random seed. Results for other models, including NeuralAtoms (Li et al., 2024), Exphormer (Shirzad et al., 2023) and GraphGPS (Rampášek et al., 2022), were reported as published in their original works.
@@ 4.3 Ablations and analyses
Long-range spoke update layer components.
Table 4: Ablation study of long-range spoke update layer components. We measure the effect of various components of ReHub on top of a GatedGCN MPNN, using the PascalVOC-SP dataset. The number of hubs used per graph (#Hubs): for 22 it is a static amount and for $\sqrt{N_{s}}$ it is dynamic per graph size. Initial hubs (Hubs Init) can be set as learned parameters or initialized from the assigned spokes as described in 3.3 where we can add a feedforward layer on the spokes (Spokes Enc) before aggregation. Reassignment is as described in 3.4. We use $k=3$ for all runs.
@@
As described in Section 3, our long-range spoke update layer is designed to handle long range communication while maintaining memory efficiency. The ablation provided in Table 4 highlights the contributions of the primary design choices, evaluated on PascalVOC-SP which contains an average of 479.4 nodes per graph. (1) Initialization of hub features from spokes vs. learned hub features as parameters, where the latter is analogous to the initialization process of Neural Atoms. The initialization scheme, outlined in Section 3.3, is compared to a configuration where hub features are learned parameters, i.e, $\bm{h}_{i_{h}}^{0}={\bm{P}}_{i_{h}}\in\mathbb{R}^{d}$. Initializing the hubs from spokes significantly improves performance, while using learned hubs results in reduced performance compared to the GNN baseline. (2) The use of a fixed vs. dynamic number of hubs. For the fixed case, we set the number of hubs to 22, which is approximately the square root of the average number of nodes, $\sqrt{479.4}\approx 21.9$. This is compared to setting the number of hubs as $N_{h}=\sqrt{N_{s}}$ dynamically, according to graph size. (3) Reassigning spokes to hubs at each layer vs. keeping the same connections fixed across the layers, is shown to improve performance. (4) Including an encoding layer for the spokes prior to aggregation further improves performance.
Table 5: Ablation study of spoke-hub assignment strategies. We analyze the impact of various strategies for hub clustering, initial hub assignment, and hub reassignment on ReHub’s performance using the PascalVOC-SP dataset. Accordingly, the results are grouped into three categories, with ReHub’s performance reported at the bottom. The clustering methods include random, balanced random, and METIS. For hub assignment and reassignment, we evaluate random, balanced random, feature similarity-based assignment, and no reassignment. Reassignment based on Attention Scores is as described in Section 3.4. We use $k=3$ for all runs. Our findings show that neither of the random strategies improve the results and actually degrade performance. This further supports our contention that our feature similarity based strategy leverages meaningful spoke-hub relationships.
Clustering Hubs Assignment Reassignment PascalVOC-SP (F1 $\uparrow$)
Random Features Similarity Attention Scores 0.3503 $\pm$ 0.0143
Balanced Random Features Similarity Attention Scores 0.3622 $\pm$ 0.0060
METIS Features Similarity Attention Scores 0.3860 $\pm$ 0.0172
METIS Random Attention Scores 0.3619 $\pm$ 0.0083
METIS Balanced Random Attention Scores 0.3618 $\pm$ 0.0126
METIS Features Similarity Attention Scores 0.3860 $\pm$ 0.0172
METIS Features Similarity No Reassignment 0.3775 $\pm$ 0.0040
METIS Features Similarity Random 0.3514 $\pm$ 0.0121
METIS Features Similarity Balanced Random 0.3486 $\pm$ 0.0106
METIS Features Similarity Attention Scores 0.3860 $\pm$ 0.0172
Assignment of spokes and hubs.
The clustering into hubs and assignment to spokes are crucial components of ReHub’s architecture. The ablation in Table 5 illustrates how different clustering and assignment strategies affect ReHub, by evaluating on PascalVOC-SP. We identify three key constraints that any assignment strategy should satisfy: (1) each spoke must be assigned to exactly $k$ hubs, (2) no hub should appear more than once in a spoke’s assignments, and (3) the number of spokes assigned to each hub should be approximately balanced. We therefore devise two baseline strategies to compare against our feature similarity based strategy: Random and Balanced Random. The Random strategy assigns to each spoke $k$ hubs sampled uniformly at random without replacement. While this ensures each spoke has exactly $k$ unique hubs, it does not guarantee a balanced distribution of spokes across hubs, and only approaches balance asymptotically for very large graphs. To address this, we propose the Balanced Random strategy, which enforces balance while maintaining some degree of randomness. This assignment task inherently involves a trade-off between computational complexity, randomness, and balance: achieving perfect balance with full randomness is computationally expensive, while a fully random assignment is efficient but unbalanced. The Balanced Random strategy trades off some randomness to achieve approximate balance with linear complexity. Specifically, we begin by randomly permuting the hub indices $0,\ldots,N_{h}-1$, and for each spoke with index $i$, assign the $k$ consecutive hubs at positions $(i\cdot u+j)\bmod N_{h}$ for $j=0,\ldots,k-1$, where $u$ is an integer coprime with $N_{h}$. This construction preserves partial randomness, guarantees $k$ unique hubs per spoke, and distributes spokes evenly across hubs. For implementation details, see Appendix A.3.
We evaluate both random assignment strategies in three modules: (1) Clustering, as defined in Section 3.3, is used for hub feature aggregation and can be performed randomly or in a balanced random manner, both following the same algorithmic logic described above with $k=1$. This is compared against METIS, which is the clustering strategy used by ReHub; results are shown in rows 1-3 of Table 5. (2) Initial Hub Assignment is evaluated on random and balanced random strategies, and also compared against ReHub’s strategy that assigns hubs based on feature similarity. Results are shown in rows 4-6 of Table 5. (3) Reassignment can be either disabled, performed randomly, balanced randomly, or with the attention score based method as described in Section 3.4. Results are shown in rows 7-10 of Table 5.
Our findings show that both random strategies degrade performance, further supporting the contention that our feature similarity based strategy leverages meaningful spoke-hub relationships.
Figure 3: Hub utilization. Histogram of used hub percentages across graphs in the PascalVOC-SP validation set, for different configurations of hub ratio $r$ and number of connected hubs $k$. Each subplot shows hub utilization across four layers. Percentages are grouped using bins of size 10%. Nearly all hubs are utilized across all layers, with the peak typically near 90%-100% usage.
Hub utilization.
@@
To gain insights into the reassignment dynamics, we measure the level of hub utilization. We define hub utilization $U$ as the number of hubs that have at least one spoke connected to them. Formally, $U=|\{i_{h}\mid{\bm{E}}_{:,i_{h}}\cdot{\bm{1}}_{N_{h}}\geq 1\}|$ where ${\bm{E}}_{:,i_{h}}$ represents the set of all spokes connected to hub $i_{h}$. The percentage of used hubs is then given by $U/N_{h}$, which reflects the proportion of hubs that are connected to at least one spoke. Figure 3 shows histograms of hub utilization across graphs, where the X-axis represents the percentage of used hubs (in bins of size 10), and the Y-axis represents the percentage of graphs falling into each bin. Separate histograms are shown for each layer (Layers 1–4) and for different configurations of hub ratio $r$ and connected hubs $k$, based on the validation split of the PascalVOC-SP dataset. The results demonstrate that in the vast majority of graphs, nearly all hubs are utilized across all layers, with the peak typically near 90%-100% usage. This finding suggests that the network maintains robust information flow between spokes and hubs, regardless of the number of hubs or the number of spokes per hub. We also present an additional metric based on the Bhattacharyya coefficient in Appendix A.4.
@@ 5 Conclusion and future work
Future work  While consistently improving computational efficiency, our spoke-hub communication maintained strong performance, establishing a foundation for future works. Additionally, our current design lacks inherent support for positional information available in geometric graphs. In future work, we aim to extend our method to support tasks requiring long-range communication on geometric graphs by incorporating positional information into the spoke-hub attention and reassignment mechanisms. We also plan to make the reassignment module learnable, and optimize it towards the prediction task to further boost accuracy. Finally, integrating ReHub into architectures like Exphormer to enable the reassignment of expander graph edges between layers presents a promising direction to further enhance long-range communication.